# Contrastive Semantic Projection: Faithful Neuron Labeling with Contrastive Examples

### 24 Apr 2026 v1

@@ 4 Experiments
We evaluate two families of internal units: (i) convolutional features from ResNet-50 and ResNet-101 [19] (channels after the 4th residual block with spatial average pooling), and (ii) Sparse Autoencoder (SAE) features with ReLU or Top-$K$ sparsity [4, 11, 14] trained on post-residual-stream representations of CLIP models (see Section˜0.A.1 for SAE details).
@@ 4.1 Improving Candidate Labels Through Contrastive Examples
We first evaluate the quality of candidate labels generated for each neuron, independently of any downstream labeling pipeline. For each neuron, we prompt the VLMs using either (i) positive‑only samples or (ii) our contrastive prompting approach. Contrastive examples are extracted according to Eq.˜1 using CLIP ViT‑B/16 on the MS COCO 2017 train set. Each prompt receives a $\times3$ grid of images (positive or contrastive; see Fig.˜3c), and each VLM provides three candidate labels. Full prompt templates are listed in Section˜0.A.2.
@@
Our analysis focuses on a CLIP ViT‑B/16 model trained on DataComp-XL [10]. We study latent representations after the final transformer block and apply an SAE trained on MS COCO 2017 (details in Section˜0.A.1). We evaluate the 200 neurons with the highest average activation magnitude to ensure that only consistently active and semantically meaningful neurons are included.
@@ 4.2 Evaluating Label Faithfulness of Contrastive Explanations
We use the ImageNet-1K validation set for probing and ImageNet-21k [24] class names as the baseline label vocabulary. For each model, we evaluate 200 units selected by highest mean activation over the probe set. For each unit, we form a positive set from the top-$30$ activating images and construct contrastives by pairing each positive image with its nearest neighbor (in CLIP ViT-B/32 embedding space) that activates the unit below its mean. We vary the amount of positive and contrastive pairs and study its effect in Section˜0.A.4.1.
We include two variants of CSP, corresponding to different values of the contrastive scaling parameter $\gamma$ as defined in Section˜3.2. Specifically, we include the default formulation with $\gamma=1$ and a less invasive variant with $\gamma=0.5$. We vary $\gamma$ and report its impact on labeling faithfulness in Section˜0.A.4.2. We additionally test against two baselines that make use of contrastive examples in Section˜0.A.4.3.
@@
We also observe similar but weaker trends for AUC and SCS (averaged over target models and reported in Table˜2, full per-model results available in Section˜0.A.3). For AUC, CSP shares the top performance with LE in the no augmentation and positive-only augmentation settings. When augmenting with contrastive labels, the gap widens slightly with CSP exceeding LE by +0.02+0.02, suggesting that CSP benefits more from contrastive labels on this metric. For SCS, however, LE consistently achieves the highest scores across all augmentation regimes, with CSP ranking second. Since LE trains linear classifiers across the full activation range using soft labels, it is naturally aligned with the SCS objective and tends to yield higher correlations. Across all labeling pipelines, the improvements from contrastive augmentation are small.
@@
Lastly, neuron #869 (ResNet101) activates for nature scenery images that include multiple trees. Contrastive images show natural scenes as well but are more varied (e.g., grass only, mountains, and snow). In the positive-only augmentation setting, CSP predicts cedar_of_lebanon, a specific tree type, which activates the neuron lower than alternatives. With the contrastive labeling augmentation, the label natural_forest_setting is introduced and gets predicted only by CSP, yielding the highest activation on the neuron. Other pipelines still assign generic labels such as outdoor_scene. We provide more qualitative examples in Section˜0.A.6.
@@ 4.3 Medical Use Case: Skin Lesion Labeling
Unlike the natural-image setting, generative models are unavailable, making the SCS metric the only faithfulness metric we report (in Table˜3) in this experiment. We use DermLIP [29], a vision-language model for dermatology, as the simulator. Results across different values of $\gamma$ in this use case are reported in Section˜0.A.5.1.
@@ Acknowledgements.
Appendix 0.A Appendix
0.A.1 SAE Training Details
@@ 0.A.2 Prompt Templates
0.A.3 Per-Model and Per-Layer Tables (Natural Images)
In Section˜4.2, we report AUC and SCS metrics averaged over the four target models. In Tables˜4 and 5, we report both of these metrics for each target model individually. Across all target models and all VLM label augmentation regimes, CSP consistently tops the AUC scores, while LE gets the highest SCS scores followed by CSP.
Table 4: AUC scores across architectures and labeling pipelines. Values are reported as mean $\pm$ standard error of the mean across the evaluated neurons and the three label augmentation regimes. The number of neurons using augmented labels (when applicable) included in parenthesis. Best-performing pipelines are highlighted in bold.
@@ 0.A.3 Per-Model and Per-Layer Tables (Natural Images)
Table 5: SCS scores across architectures and labeling pipelines. Values are reported as mean $\pm$ standard error of the mean (in %) across the evaluated neurons and the three label augmentation regimes. The number of neurons using augmented labels (when applicable) included in parenthesis. Best-performing pipelines are highlighted in bold.
@@
0.A.4 Ablation Results (Natural Images)
0.A.4.1 Amount of Samples $k$
Table 6: DMA scores for labeling pipelines using baseline labels only, under different values $k$ of positive and contrastive samples. Only CSP makes use of the contrastive samples. LE is $k$-agnostic as it relies on the entire activation range and is therefore reported as a separate row. Values are reported as mean $\pm$ standard error of the mean (in %). For each $k$ value, the best-performing pipeline is highlighted in bold.
@@ 0.A.4.1 Amount of Samples $k$
We vary the number of positive and contrastive samples, $k\in\{10,20,\dots,100\}$, and report the resulting DMA scores in Table˜6 for ResNet50 and SAE-TopK. Since LE uses the full activation range, it is unaffected by $k$ and is reported separately. Across both models, the best results are achieved by CSP variants, which typically peak at intermediate values around k=30k=304040. In contrast, CLIP-Dissect and SemanticLens perform best at smaller values, usually $k=10$ or $20$.
@@
0.A.4.2 Subtraction Factor $\gamma$
Table 7: DMA scores for CSP using baseline labels only, under different $\gamma$ values, for ResNet50 and SAE-TopK. Values are reported as mean $\pm$ (in %) standard error of the mean. Best variant per architecture is in bold.
@@ 0.A.4.2 Subtraction Factor $\gamma$
We study CSP for $\gamma\in\{0,0.1,\dots,1.0\}$ without label augmentation. Table˜7 reports the resulting DMA scores for ResNet50 and SAE-TopK. For ResNet50, performance improves steadily as $\gamma$ increases, with the best result at $\gamma=1.0$. SAE-TopK shows a similar overall trend, but peaks at $\gamma=0.9$, with a small drop at $\gamma=1.0$. Overall, the results indicate that stronger subtraction is generally beneficial, while full projection can be slightly suboptimal in some cases.
0.A.4.3 Contrastive Baselines
Table 8: DMA, AUC, and SCS scores without label augmentation, comparing vanilla CSP, CSP with $\gamma=0.5$, a contrastive score-diff baseline (see Eq.˜18), and a contrastive embedding-diff baseline (see Eq.˜19), for ResNet50 and SAE-TopK. Values are reported as mean $\pm$ standard error of the mean. Best variant per metric and architecture is shown in bold.
@@ 0.A.4.3 Contrastive Baselines
Table˜8 reports results without label augmentation for ResNet50 and SAE-TopK. For ResNet50, vanilla CSP performs best on all three metrics, improving over the score-difference baseline by $16.34\%$ in DMA, $7.41\%$ in AUC, and $8.39\%$ in SCS, and over the embedding-difference baseline by $21.26\%$, $8.75\%$, and $11.99\%$, respectively.
@@
0.A.5 Ablation Results (ISIC)
0.A.5.1 Subtraction Factor $\gamma$
Table 9: SCS scores for CSP under different $\gamma$ values on the ISIC use case using DermLIP as the simulator. Values are reported as mean $\pm$ (in %) standard error of the mean. Best variant is in bold.
@@ 0.A.5.1 Subtraction Factor $\gamma$
Table˜9 reports SCS scores for CSP with $\gamma\in\{0.1,0.2,\dots,1.0\}$. In both the positive-only and contrastive augmentation settings, performance peaks at an intermediate value of $\gamma$ ($0.4$ and $0.5$, respectively) and declines as $\gamma$ approaches full projection. This again highlights the importance of tuning $\gamma$ for CSP. Unlike the natural-image setting, the best performance in this medical use case is achieved with partial rather than near-complete projection.
0.A.6 Additional Qualitative Results (Natural Images)
We provide additional qualitative examples for the natural images experiment in Section˜4.2. Examples without label augmentation are shown in Figs.˜7, 8, 9 and 10. Examples comparing labels produced under the positive-only and contrastive augmentation strategies are shown in Figs.˜11, 12, 13 and 14. Additional cases illustrating documented over-projection are presented in Fig.˜15.
Figure 7: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure 8: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure 9: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure 10: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure 11: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure 12: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure 13: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure 14: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure 15: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset. These examples showcase cases where CSP fails due to over-projection.

### 29 Apr 2026 v2

@@ Abstract
footnotetext: Code and experiment scripts are available at https://github.com/oussamabouanani/neurolens
@@ 4 Experiments
We evaluate two families of internal units: (i) convolutional features from ResNet-50 and ResNet-101 [19] (channels after the 4th residual block with spatial average pooling), and (ii) Sparse Autoencoder (SAE) features with ReLU or Top-$K$ sparsity [4, 11, 14] trained on post-residual-stream representations of CLIP models (see Section˜A.1 for SAE details).
@@ 4.1 Improving Candidate Labels Through Contrastive Examples
We first evaluate the quality of candidate labels generated for each neuron, independently of any downstream labeling pipeline. For each neuron, we prompt the VLMs using either (i) positive‑only samples or (ii) our contrastive prompting approach. Contrastive examples are extracted according to Eq.˜1 using CLIP ViT‑B/16 on the MS COCO 2017 train set. Each prompt receives a $\times3$ grid of images (positive or contrastive; see Fig.˜3c), and each VLM provides three candidate labels. Full prompt templates are listed in Section˜A.2.
@@
Our analysis focuses on a CLIP ViT‑B/16 model trained on DataComp-XL [10]. We study latent representations after the final transformer block and apply an SAE trained on MS COCO 2017 (details in Section˜A.1). We evaluate the 200 neurons with the highest average activation magnitude to ensure that only consistently active and semantically meaningful neurons are included.
@@ 4.2 Evaluating Label Faithfulness of Contrastive Explanations
We use the ImageNet-1K validation set for probing and ImageNet-21k [24] class names as the baseline label vocabulary. For each model, we evaluate 200 units selected by highest mean activation over the probe set. For each unit, we form a positive set from the top-$30$ activating images and construct contrastives by pairing each positive image with its nearest neighbor (in CLIP ViT-B/32 embedding space) that activates the unit below its mean. We vary the amount of positive and contrastive pairs and study its effect in Section˜A.4.1.
We include two variants of CSP, corresponding to different values of the contrastive scaling parameter $\gamma$ as defined in Section˜3.2. Specifically, we include the default formulation with $\gamma=1$ and a less invasive variant with $\gamma=0.5$. We vary $\gamma$ and report its impact on labeling faithfulness in Section˜A.4.2. We additionally test against two baselines that make use of contrastive examples in Section˜A.4.3.
@@
We also observe similar but weaker trends for AUC and SCS (averaged over target models and reported in Table˜2, full per-model results available in Section˜A.3). For AUC, CSP shares the top performance with LE in the no augmentation and positive-only augmentation settings. When augmenting with contrastive labels, the gap widens slightly with CSP exceeding LE by +0.02+0.02, suggesting that CSP benefits more from contrastive labels on this metric. For SCS, however, LE consistently achieves the highest scores across all augmentation regimes, with CSP ranking second. Since LE trains linear classifiers across the full activation range using soft labels, it is naturally aligned with the SCS objective and tends to yield higher correlations. Across all labeling pipelines, the improvements from contrastive augmentation are small.
@@
Lastly, neuron #869 (ResNet101) activates for nature scenery images that include multiple trees. Contrastive images show natural scenes as well but are more varied (e.g., grass only, mountains, and snow). In the positive-only augmentation setting, CSP predicts cedar_of_lebanon, a specific tree type, which activates the neuron lower than alternatives. With the contrastive labeling augmentation, the label natural_forest_setting is introduced and gets predicted only by CSP, yielding the highest activation on the neuron. Other pipelines still assign generic labels such as outdoor_scene. We provide more qualitative examples in Section˜A.6.
@@ 4.3 Medical Use Case: Skin Lesion Labeling
Unlike the natural-image setting, generative models are unavailable, making the SCS metric the only faithfulness metric we report (in Table˜3) in this experiment. We use DermLIP [29], a vision-language model for dermatology, as the simulator. Results across different values of $\gamma$ in this use case are reported in Section˜A.5.1.
@@ Acknowledgements.
Appendix A Appendix
A.1 SAE Training Details
@@ A.2 Prompt Templates
A.3 Per-Model and Per-Layer Tables (Natural Images)
In Section˜4.2, we report AUC and SCS metrics averaged over the four target models. In Tables˜A.1 and A.2, we report both of these metrics for each target model individually. Across all target models and all VLM label augmentation regimes, CSP consistently tops the AUC scores, while LE gets the highest SCS scores followed by CSP.
Table A.1: AUC scores across architectures and labeling pipelines. Values are reported as mean $\pm$ standard error of the mean across the evaluated neurons and the three label augmentation regimes. The number of neurons using augmented labels (when applicable) included in parenthesis. Best-performing pipelines are highlighted in bold.
@@ A.3 Per-Model and Per-Layer Tables (Natural Images)
Table A.2: SCS scores across architectures and labeling pipelines. Values are reported as mean $\pm$ standard error of the mean (in %) across the evaluated neurons and the three label augmentation regimes. The number of neurons using augmented labels (when applicable) included in parenthesis. Best-performing pipelines are highlighted in bold.
@@
A.4 Ablation Results (Natural Images)
A.4.1 Amount of Samples $k$
Table A.3: DMA scores for labeling pipelines using baseline labels only, under different values $k$ of positive and contrastive samples. Only CSP makes use of the contrastive samples. LE is $k$-agnostic as it relies on the entire activation range and is therefore reported as a separate row. Values are reported as mean $\pm$ standard error of the mean (in %). For each $k$ value, the best-performing pipeline is highlighted in bold.
@@ A.4.1 Amount of Samples $k$
We vary the number of positive and contrastive samples, $k\in\{10,20,\dots,100\}$, and report the resulting DMA scores in Table˜A.3 for ResNet50 and SAE-TopK. Since LE uses the full activation range, it is unaffected by $k$ and is reported separately. Across both models, the best results are achieved by CSP variants, which typically peak at intermediate values around k=30k=304040. In contrast, CLIP-Dissect and SemanticLens perform best at smaller values, usually $k=10$ or $20$.
@@
A.4.2 Subtraction Factor $\gamma$
Table A.4: DMA scores for CSP using baseline labels only, under different $\gamma$ values, for ResNet50 and SAE-TopK. Values are reported as mean $\pm$ (in %) standard error of the mean. Best variant per architecture is in bold.
@@ A.4.2 Subtraction Factor $\gamma$
We study CSP for $\gamma\in\{0,0.1,\dots,1.0\}$ without label augmentation. Table˜A.4 reports the resulting DMA scores for ResNet50 and SAE-TopK. For ResNet50, performance improves steadily as $\gamma$ increases, with the best result at $\gamma=1.0$. SAE-TopK shows a similar overall trend, but peaks at $\gamma=0.9$, with a small drop at $\gamma=1.0$. Overall, the results indicate that stronger subtraction is generally beneficial, while full projection can be slightly suboptimal in some cases.
A.4.3 Contrastive Baselines
Table A.5: DMA, AUC, and SCS scores without label augmentation, comparing vanilla CSP, CSP with $\gamma=0.5$, a contrastive score-diff baseline (see Eq.˜18), and a contrastive embedding-diff baseline (see Eq.˜19), for ResNet50 and SAE-TopK. Values are reported as mean $\pm$ standard error of the mean. Best variant per metric and architecture is shown in bold.
@@ A.4.3 Contrastive Baselines
Table˜A.5 reports results without label augmentation for ResNet50 and SAE-TopK. For ResNet50, vanilla CSP performs best on all three metrics, improving over the score-difference baseline by $16.34\%$ in DMA, $7.41\%$ in AUC, and $8.39\%$ in SCS, and over the embedding-difference baseline by $21.26\%$, $8.75\%$, and $11.99\%$, respectively.
@@
A.5 Ablation Results (ISIC)
A.5.1 Subtraction Factor $\gamma$
Table A.6: SCS scores for CSP under different $\gamma$ values on the ISIC use case using DermLIP as the simulator. Values are reported as mean $\pm$ (in %) standard error of the mean. Best variant is in bold.
@@ A.5.1 Subtraction Factor $\gamma$
Table˜A.6 reports SCS scores for CSP with $\gamma\in\{0.1,0.2,\dots,1.0\}$. In both the positive-only and contrastive augmentation settings, performance peaks at an intermediate value of $\gamma$ ($0.4$ and $0.5$, respectively) and declines as $\gamma$ approaches full projection. This again highlights the importance of tuning $\gamma$ for CSP. Unlike the natural-image setting, the best performance in this medical use case is achieved with partial rather than near-complete projection.
A.6 Additional Qualitative Results (Natural Images)
We provide additional qualitative examples for the natural images experiment in Section˜4.2. Examples without label augmentation are shown in Figs.˜A.1, A.2, A.3 and A.4. Examples comparing labels produced under the positive-only and contrastive augmentation strategies are shown in Figs.˜A.5, A.6, A.7 and A.8. Additional cases illustrating documented over-projection are presented in Fig.˜A.9.
Figure A.1: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure A.2: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure A.3: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure A.4: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset.
Figure A.5: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure A.6: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure A.7: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure A.8: Examples of improving labeling faithfulness (via DMA score) after augmenting the labels with contrastive images.
Figure A.9: Examples of labels predicted by different labeling pipelines and their corresponding DMA scores, using only the baseline labels dataset. These examples showcase cases where CSP fails due to over-projection.